### v1

# Progress-Aware Video Frame Captioning

Abstract
While image captioning provides isolated descriptions for individual images, and video captioning offers one single narrative for an entire video clip, our work explores an important middle ground: progress-aware video captioning at the frame level. This novel task aims to generate temporally fine-grained captions that not only accurately describe each frame but also capture the subtle progression of actions throughout a video sequence. Despite the strong capabilities of existing leading vision language models, they often struggle to discern the nuances of frame-wise differences. To address this, we propose ProgressCaptioner, a captioning model designed to capture the fine-grained temporal dynamics within an action sequence. Alongside, we develop the FrameCap dataset to support training and the FrameCapEval benchmark to assess caption quality. The results demonstrate that ProgressCaptioner significantly surpasses leading captioning models, producing precise captions that accurately capture action progression and set a new standard for temporal precision in video captioning. Finally, we showcase practical applications of our approach, specifically in aiding keyframe selection and advancing video understanding, highlighting its broad utility.  [^1] Project webpage: https://vision.cs.utexas.edu/projects/ProgressCaptioner.
Work conducted as an independent researcher.
1 Introduction
Visual captioning [38]—the task of generating textual descriptions of visual content—is a fundamental problem in computer vision with extensive practical applications. Existing captioning paradigms are broadly divided into two categories: image captioning and video captioning, with a clear distinction between them. Image captioning [23] generates a single, isolated description for each image, with no contextual linkage among different images. In contrast, video captioning [1] assigns a single caption for the entire video clip, aggregating information across frames without addressing the specifics of individual frames.
Figure 1: We propose progress-aware video frame captioning (bottom), which aims to generate a sequence of captions that capture the temporal dynamics within a video. Unlike traditional image and video captioning (top) that focus on broad event-level descriptions, our task delves into the detailed, progressive dynamics of an action, necessitating precise, temporally fine-grained capabilities. Blue text highlights how the progress-aware captions build successively on the earlier content to highlight what is changing.
Figure 1 illustrates this dichotomy. Employing an image captioning model like BLIP [37] to describe each frame of the video results in captions that are local, not temporally context-aware, and may exhibit little variation across the sequence. Conversely, video captioning provides a global, not temporally fine-grained overview of the event, as exemplified by the YouCook2 [84] ground truth label “scramble the eggs in the wok”. In both scenarios, the nuances of how the action unfolds over time are missed. This raises the question: Can we develop temporally fine-grained captions that capture the subtle, progressive nature of action sequences? Figure 1 (bottom) illustrates what we seek.
Having such progress-aware captions could benefit a great variety of downstream tasks, bringing improved video understanding [78, 73], more precise video retrieval [67, 68, 66], and enriched video generation [58, 50]. Moreover, such a capability could open up new AR/VR and robotics applications. For instance, in AI coaching, a captioning system could meticulously analyze an expert’s tennis forehand, simplifying the learning process for users. Similarly, for how-to video creation, it could elicit and describe the key object state changes at each stage (e.g., “how to make whipped cream”)—useful for both content creators as well as visually-impaired users learning a new skill.
Towards this end, we introduce a novel captioning taskprogress-aware video captioning at the frame level. This task involves generating captions that not only coherently depict action progression but also tailor each description specifically to its corresponding frame. [^2] Without loss of generality, we obtain the input frame sequence by uniformly sampling from the action clips at a fixed rate (1FPS). These frames may or may not demonstrate visual action progression from one to the next, demanding that the model discerns the difference when generating progress-aware captions. Our task is uniquely characterized by its demand for fine-grained temporal sensitivity. By “fine-grained”, we refer to generating detailed, frame-level descriptions that elucidate the stages or procedural steps of the action, effectively conveying how the action is performed throughout the video sequence.
As discussed, traditional video captioning [9, 70, 85, 63] settles for broad event-level descriptions, where a description like scrambling eggs for the video in Figure 1 would be considered entirely accurate. In contrast, we seek progress-aware captions that detail each stage of the action, such as “eggs still intact”, “starting to cook” and “fully mixed”. While recent works [8, 61, 11, 82, 6, 12] enhance the overall descriptiveness of video captions, they continue to produce a single video-level description without distinguishing the nuances between individual frames. Our task delves deeper, exploring how each frame contributes to the narrative of the action’s progression, thereby setting a new standard for fine-grained temporal precision in video captioning.
Figure 2: Issues of existing VLMs in video frame captioning: (1) Lack of temporal granularity. See captions for frames 2 and 3, produced by Gemini-1.5-Pro (row 2), which fail to distinguish subtle differences between the frames. (2) Temporal hallucination. See frame 2’s caption produced by GPT-4o (row 1), which inaccurately suggests progression that is not visible.
Despite great advancements of vision language models (VLMs) [37, 75, 80, 42, 41, 40, 62, 2, 33, 82] that have markedly improved visual captioning, we observe that these models still struggle with this nuanced task. Two main issues persist: first, the lack of temporally fine-grained captions; when shown adjacent frames that depict subtle variations in action progression (such as frames 2 and 3 in Figure 2), the generated captions can be overly coarse, failing to differentiate between the frames (see row 2, Gemini-1.5-Pro’s captions). Second, we identify and term a notable issue of “temporal hallucination”, where the captions suggest temporal progression in disagreement with what the visual frames exhibit. See frame 2 of Figure 2, where GPT-4o’s generated captions (row 1) incorrectly advance the action sequence. The prevalence of such errors can be attributed to models’ reliance on the common statistics of activity sequences, which mistakenly override matching specific statements to specific frames. Meanwhile, image captioners—even if trained with fine-grained annotations—treat frames in isolation and hence lack the temporal context to say what is progressing versus what is present.
We propose ProgressCaptioner, a captioning model designed to generate progress-aware frame-level captions. This model is developed through an innovative strategy that interleaves pseudo labeling with two learning stages. In the first stage, we employ multiple VLMs to generate consensus pseudo captions for pairs of frames, and we devise two novel evaluation tasks to automatically assess their quality. In the second stage, armed with that initial model, we apply a two-frame sliding window to generate more precise captions across the full frame sequence, and we generalize our custom evaluation tasks to identify high-quality caption sequences. The final model, ProgressCaptioner, is then trained on this comprehensive dataset, which we term FrameCap, to excel in producing temporally fine-grained captions that accurately characterize action progress.
To assess the quality of frame-wise captions and benchmark ProgressCaptioner against leading VLMs, we introduce the FrameCapEval benchmark comprised of videos from four public video action datasets. ProgressCaptioner consistently outperforms leading open-source VLMs with a 1.8$\times$ to 2.7$\times$ improvement in caption quality and also achieves the highest selection rate in user studies, even surpassing the much larger proprietary models GPT-4o [2] and Gemini-1.5-Pro [53]. Finally, we highlight potential applications enabled by our advanced captions: keyframe selection and enhanced video understanding. We hope that our task, model, and benchmark can inspire future development in temporally fine-grained video captioning.
2 Related Work
Image Captioning Image captioning has been extensively studied in recent years [13, 76, 60, 3]. A related line of work is image difference captioning, where the task is to describe differences between two images [27, 48, 56] or sets of images [15]. Building on the success of generative models, recent benchmarks [4, 28, 77] challenge models to distinguish between two visually similar images, advancing fine-grained image understanding. However, all the above models are restricted to static (typically synthesized) image pairs and address coarse-grained differences like object presence or absence. Temporal intricacies—accurately describing how an action progresses—remain unexplored.
Video Captioning Video captioning [1] aims to produce a single description that encapsulates a video clip. While traditional benchmarks [9, 70, 85, 63] offer a brief one-sentence caption for each video, recent efforts expand this scope, extending captioning to hours-long videos [26], enriching the granularity of details [11, 8, 61], enhancing caption uniqueness [49], integrating a casual temporal narrative [46], or introducing LLM summarization [34].
Adjacent to traditional video captioning, are the tasks of visual storytelling [25, 36] (creating a coherent story for a sequence of snapshots), dense video captioning [30, 74, 86] (temporal localization and captioning of all events in an untrimmed video), and audio description [20, 21, 22] (detailed narrations of visual events in videos (e.g., movies) for visually impaired audiences). However, all these works still address “what is happening” at a coarse-grained event level, e.g., noting that someone is making a souffle within a specific time range. The ability to break down frame-level details—such as whisking egg whites, folding ingredients, and observing the souffle rising—is still lacking.
Vision Language Models Recent advancements in VLMs [37, 75, 80, 42, 41, 40, 62, 2, 33, 82] have greatly enhanced the capabilities of both image and video captioning. Despite their strong performance, VLMs often exhibit “hallucination” [65, 19], and preference learning [52, 83] has proven effective in mitigating this issue.
Compared to image-LMs, video-LMs crucially require the integration of temporal dynamics understanding, spurring a series of work on evaluating temporal perception [18, 69, 43, 39]. While these assessments ensure that a model can generate an accurate overall video summary or answer general questions, they entail neither temporal localization nor discernment of fine-grained differences between frames. The OSCaR benchmark [47] focuses on object state change (OSC) captioning, yet it is limited to just three frames and specifically OSC videos, with models and captions not publicly released yet preventing direct comparison. Additionally, their approach relies on human annotation and a single advanced GPT model. In contrast, our approach features a scalable data collection pipeline that reduces reliance on these labor-intensive resources, employs novel automatic evaluation tasks, and broadens the scope beyond OSC videos. Finally, unlike methods for long-form video and event localization with VLMs [44, 10, 54], our focus is distinctly more temporally fine-grained, concentrating on how individual frames evolve within a single event.
3 Approach
We delve into the specific challenges of our progress-aware video frame captioning problem in Sec. 3.1 and outline ProgressCaptioner’s development in Sec. 3.2.
3.1 Progress-aware Video Frame Captioning
Problem Formulation Our objective is to develop a captioning model that, given a video, produces accurate and temporally fine-grained captions. Formally, for a sequence of $T$ frames, denoted as $\mathcal{V}=\{v_{i}\}_{i=1}^{T}$, the captioning model generates a corresponding sequence of captions $\mathcal{C}=\{c_{i}\}_{c=1}^{T}$, where each $c_{i}$ describes the $i$-th frame $v_{i}$. This captioning process is underpinned by three key requirements: (1) Accuracy, where each caption $c_{i}$ must faithfully represent what is visually occurring in frame $v_{i}$, without hallucinating from the context of other frames; (2) Temporal Specificity, where each caption $c_{i}$ specifically attends to $v_{i}$, without being overly generic to be applicable to multiple frames in the sequence; (3) Progressive Coherence: The sequence of captions $\{c_{i}\}_{i=1}^{T}$ should build upon each other to reflect the essential changes in the action over time.
FrameCap Dataset To train our captioning model, we require a dataset that pairs frame sequences ($\mathcal{V}$) with corresponding captions ($\mathcal{C}$). Existing datasets [9, 70, 85, 63] provide only a single, generic caption for an entire video clip, lacking the frame-wise caption format we need. To address this gap and train our model, we develop the FrameCap dataset. Given the prohibitive expense of collecting human-labeled caption sequences as our ground truth ($\mathcal{C}$), especially at scale, we leverage leading VLMs as powerful tools to create a pseudo caption sequence $\hat{\mathcal{C}}$ from $\mathcal{V}$. For video sources, we refer to two large-scale datasets that focus on fine-grained human activities: HowToChange [72] (featuring object state change videos from YouTube) and COIN [59] (featuring daily activities from YouTube).
Caption Sequence Construction Prompting VLMs for our desired caption sequence is nontrivial. We identify two key problems: (1) Input considerations: how many context frames from $\{v_{i}\}_{i}^{T}$ should be provided? (2) Output assessment: what issues arise in VLM-generated captions, and how can we filter to retain only high-quality ones? To explore these questions, we conduct preliminary experiments by prompting leading VLMs to perform the frame-wise captioning task. We share our findings below.
Observation I Intuitively, inputting all $T$ frames would seem best. However, current VLM capabilities do not support this extensive context. Specifically, providing too many frames at once often leads to descriptions that lack detail and exhibit temporal inaccuracies, with VLMs also risking memory overload, as similarly observed in [11]. Conversely, providing a single frame at a time reduces the task to image captioning, which is not optimal either, resulting, resul in captions that lack temporal context and coherence.
Figure 3: Captioning outcomes using Gemini-1.5-Pro [53].
Figure 3 shows a representative trial with Gemini-1.5-Pro [53]. Inputting the full sequence (case (a)) yields brief per-frame descriptions with temporal misalignment (i.e., the second caption erroneously describes what is visually occurring in the third frame). On the other hand, captioning frames in isolation (case (b)) removes essential temporal context, where the model mistakes the initial stage of a tennis serve for the follow-through of a forehand swing. These findings underscore the importance of finding a balanced approach and motivate us to adopt a frame pair as the stepping stone of our captioning model development.
Figure 4: Framework of ProgressCaptioner, designed in two stages. In Stage-I, we prepare frame pairs and generate corresponding caption pairs using multiple VLMs. Each pair undergoes our designed progression detection and caption matching evaluations, to decide if they are selected for model supervised fine-tuning or rejected, with the latter contributing to preference data to aid in model preference learning. The Stage-I model training then proceeds using this collected data. In Stage-II, the trained stage-I model labels frame sequences with a two-frame sliding window, in conjunction with other VLMs. These sequences are again assessed through progression detection and caption matching to classify them as selected or rejected. All collected data from both stages contribute to the final training of ProgressCaptioner.
Observation II Next, building upon the use of a frame pair ($v_{1},v_{2}$), is the caption pair ($\hat{c}_{1},\hat{c}_{2}$) produced by existing VLMs of sufficient quality to be directly adopted? Our preliminary experiments reveal two main issues: (a) lack of temporal granularity, and (b) temporal hallucination, as showcased in Figure 2. To dissect these issues, we analyze the captions in relation to the visual progression between frames $v_{1}$ and $v_{2}$. Specifically, if there is a visible progression from $v_{1}$ to $v_{2}$ (e.g., the slight peeling of a shrimp’s shell from frame 2 to frame 3 in Figure 2), the captions should adequately reflect this change. Overly similar captions in such a scenario signify a failure in temporal granularity. Conversely, when there is no change between frames (e.g. frames 1 and 2 in Figure 2), the captions should remain consistent. We deem it a temporal hallucination when captions erroneously indicate progression in disagreement with the visuals. This issue often arises in VLMs equipped with strong language decoders, which can introduce a language bias toward generating unwarranted narrative changes.
3.2 ProgressCaptioner
The observations above drive the design of our model, ProgressCaptioner, which unfolds in two stages. Based on our findings that current VLMs have trouble maintaining caption quality when handling extensive $T$-frame inputs, our approach begins with frame pair captioning. In the first stage, we develop a ProgressCaptioner to excel at describing the nuances between adjacent frames. The second stage then leverages the first-stage model to pseudo label the full $T$-frame sequence with a two-frame sliding window. This staged approach refines caption quality along with model development, enhancing the captioning process iteratively with more precise pseudo labels.
Frame Pair Data Preparation Starting with a frame pair $\mathbf{v}=(v_{1},v_{2}$), we employ $K$ captioning models to generate an initial set of caption pairs $\{(\hat{c}_{1},\hat{c}_{2})\}_{1}^{K}$. Acknowledging the potential inaccuracies in these captions, as per observation II, we design two automatic evaluation tasks to assess caption quality. The first task, progression detection, examines progress awareness: it checks whether the captions appropriately reflect visual changes between $v_{1}$ and $v_{2}$. Specifically, an LLM assesses each caption pair $(\hat{c}_{1},\hat{c}_{2})_{k}$ to determine if they suggest a visible physical change. We utilize majority voting across multiple LLMs’ assessments for all $K$ caption pairs to establish a consensus visual-change label. Caption pairs that align with this consensus are marked as passing; others are marked as failing.
For pairs passing progression detection, we proceed to our second evaluation task—caption matching—to assess how precisely $\hat{c}_{1}$ and $\hat{c}_{2}$ describe $v_{1}$ and $v_{2}$, respectively. The task is designed as a multi-choice question format, where a VLM is given $\hat{c}_{1}$, $\hat{c}_{2}$, and an “unsure” option, and tasked with matching the correct caption to each frame. A caption pair is considered good if the evaluation VLM correctly identifies $\hat{c}_{1}$ for $v_{1}$ and $\hat{c}_{2}$ for $v_{2}$. Because the captions will all be topically related, this is essentially a matching task with “hard negatives” that lets us automatically gauge the precision of the proposed captions for the target images.
This automatic pipeline distinguishes between high-quality caption pairs, denoted by $\mathbf{\hat{c}}^{+}=(\hat{c}_{1}^{+},\hat{c}_{2}^{+})$, and those that exhibit inaccuracies or hallucinations, denoted by $\mathbf{\hat{c}}^{-}=(\hat{c}_{1}^{-},\hat{c}_{2}^{-})$, forming training data for Stage I. [^3] We encourage readers to view data examples provided in Supp. for a better understanding of our data refinement process, as well as details on pseudo labeling, prompts used, and the $K$ VLMs we employ.
Stage I Training Following the success of versatile VLMs in captioning tasks [71, 42, 61, 34], we initialize ProgressCaptioner with the LLAVA-OV-7B [33] checkpoint to inherit its pretrained capabilities. Stage-I training utilizes frame and caption pair data collected on HowToChange and COIN YouTube videos <$\mathbf{v},\mathbf{\hat{c}}^{+},\mathbf{\hat{c}}^{-}$>  through two principal methods: supervised fine-tuning (SFT) and direct preference optimization (DPO). The SFT process is straightforward given our dataset; we perform instructional tuning to tailor the general capabilities of the original VLM to our specific frame-wise captioning requirements using <$\mathbf{v},\mathbf{\hat{c}}^{+}$>. The subsequent DPO step targets the prevalent issue of hallucination in VLMs and is innovatively driven by our proposed automatic evaluation critics. Preference optimization [52] in LLM training typically requires human-provided preference data to steer LLM responses towards more desirable outputs. Here, we employ progression detection and caption matching to automatically construct preference data $\mathbf{\hat{c}}^{+}$ and $\mathbf{\hat{c}}^{-}$, eliminating the reliance on manual labeling. This preference data <$\mathbf{v},\mathbf{\hat{c}}^{+},\mathbf{\hat{c}}^{-}$>  is adopted in DPO training to further enhance model performance with feedback from LLM and VLM evaluations.
Frame Sequence Data Preparation The second stage expands our pseudo labeling scheme from 2 to $T$ frames, where our Stage-I ProgressCaptioner first generates captions using a two-frame sliding window. To increase data diversity and volume, we also incorporate captions produced by other VLMs, with both two-frame and full $T$-frame contexts, since captions of low-quality are also useful (after undergoing our evaluation tasks, those that are rejected enrich the preference data). Once the initial set of caption sequences is generated, we conduct progression detection to identify $M$ visually distinct frames from the original $T$-frame sequence, denoted as $\mathcal{V}_{M}=\{v_{i}\}_{i=1}^{M}$, using majority voting; $M$ varies based on the distinctiveness of each frame sequence’s content. The caption matching task is then employed to encompass $M$ frames, with a selection pool of all $M$ captions, $\mathcal{\hat{C}}_{M}=\{\hat{c}_{i}\}_{i=1}^{M}$, plus an “unsure” option. A high-quality caption sequence $\mathcal{\hat{C}^{+}}$ is identified when the evaluation VLM correctly selects $\hat{c}_{i}$ for $v_{i}$ across all frames. Conversely, a caption sequence is deemed problematic, $\mathcal{\hat{C}^{-}}$, if the VLM incorrectly answers more than half of the caption selections. This process forms our Stage-II data.
Stage II Training Following the same pipeline as stage I, ProgressCaptioner is first trained through SFT using data prepared during both stages, which includes frame-caption pairs <$\mathbf{v},\mathbf{\hat{c}}^{+}$> and frame-caption sequences <$\mathcal{V},\mathcal{\hat{C}}^{+}$>. Subsequently, we conduct DPO with preference data collected from both stages <$\mathbf{v},\mathbf{\hat{c}}^{+},\mathbf{\hat{c}}^{-}$> and <$\mathcal{V},\mathcal{\hat{C}}^{+},\mathcal{\hat{C}}^{-}$>  to further refine model performance and mitigate hallucination. This sequential approach results in our final captioning model. The framework is illustrated in Figure 4.
4 Experiments
 Model Size HTC COIN Penn&K
Cap Prog Cap Prog Cap Prog
Proprietary models
Gemini-1.5-Pro [53] (img) - 28.4 59.7 24.3 58.6 15.3 51.2
Gemini-1.5-Pro [53] - 31.4 63.8 25.0 63.8 17.6 60.3
GPT-4o [2] - 32.4 64.2 21.3 58.4 18.2 63.2
Open-source models
Idefics2 [31] 8B 2.0 54.4 2.9 52.2 12.5 50.9
VILA [40] 8B 6.9 53.6 5.1 48.2 15.9 51.4
Qwen2-VL [62] 7B 13.7 69.6 11.0 70.8 8.5 58.8
LLAVA-Video [82] 7B 3.9 59.3 8.8 53.0 9.7 51.8
LLAVA-OV [33] (img) 7B 5.9 56.3 17.6 55.4 11.9 55.5
LLAVA-OV [33] 7B 7.8 59.0 5.9 57.3 5.1 50.8
ProgressCaptioner (ours) 7B 37.3 73.6 32.3 66.1 31.3 63.7
Table 1: Results on the FrameCapEval Benchmark, composed of video from four public datasets. Cap and Prog denote caption matching and progression detection accuracy, respectively. ProgressCaptioner greatly outperforms SOTA open-source VLMs and even the leading proprietary models, despite being a 7B model. The best results are bolded and underlined, the second best are bolded, and the third best are italicized. Moreover, the results confirm our model’s generalizability from in-domain datasets (HTC for HowToChange and COIN) to external datasets not seen during training (Penn&K, representing Penn Action and Kinetics).
Figure 5: Qualitative comparisons of ProgressCaptioner with SOTA VLMs on three action sequences. For sequences (a) and (b), only the middle frame predictions are displayed. See Supp. for all models’ predictions on full sequences and more examples. Inaccuracies in descriptions are highlighted in red. Even top VLMs often produce descriptions that misalign with the corresponding frames, while ProgressCaptioner delivers hallucination-free and temporally fine-grained captions, including phrases explicitly calling out progress (blue).
Figure 6: User study results comparing ProgressCaptioner with top competitors show it as the most preferred model (see text).
Figure 7: ProgressCaptioner facilitates keyframe selection and enriches the selected keyframes with progress-aware descriptions.
We tackle two questions below: (1) How to evaluate frame-wise caption quality of existing models? And how does ProgressCaptioner perform? (Sec. 4.1); (2) What applications are enabled by precise progress-aware captions? (Sec. 4.2)
4.1 Benchmarking Video Frame Captioning
To systematically evaluate the quality of generated video frame-wise captions, we propose the FrameCapEval benchmark. Our benchmark sets a new standard for temporal precision in video captioning, assessing captions based on the criteria established in Sec. 3.1.
Evaluation Metrics We employ the automatic evaluation tasks of progression detection and caption matching (Sec. 3.2), reporting accuracy with Llama-3.1-70B-Instruct [14] as the evaluation LLM and Gemini-1.5-Pro [53] as the evaluation VLM (see Supp. for details). Additionally, we enhance our evaluation with a user study of 15 participants, reporting the average selection rate.
Benchmark Data Curation We establish the FrameCapEval benchmark, featuring videos from four action understanding datasets: HowToChange [72] and COIN [59] (on which ProgressCaptioner was trained), along with Penn Action [81] and Kinetics [7], which are unseen in training and serve to assess generalization capabilities. We are mindful of the single frame bias [32] and manually verify all videos to exclude sequences lacking fine-grained action progression. This process yields a final set of 684 videos.
Baselines We evaluate an array of state-of-the-art VLMs, including two proprietary models, GPT-4o [2] and Gemini-1.5-Pro [53], and five open-source models—Idefics2 [31], VILA [40], Qwen2-VL [62], LLAVA-Video [82], and LLAVA-OV [33], all of which process the entire $T$-frame sequences as input. To ensure a comprehensive evaluation, we also incorporate image captioning baselines using Gemini-1.5-Pro and LLAVA-OV. We select open-source VLM variants with fewer than 10B parameters for computational efficiency and a fair comparison with our ProgressCaptioner, which is an 7B model. The closed proprietary models are much larger and trained with much more extensive data; we include them as a useful reference point, but stress that they do not constitute an apples-to-apples comparison, to the disadvantage of our ProgressCaptioner.
Implementation ProgressCaptioner is constructed with SigLIP [79] as the vision encoder and Qwen2 [62] as the language model, linked through a projector, and initialized from the LLAVA-OV-7B checkpoint [33]. For final-stage model training, we finetune ProgressCaptioner on 274.5K data samples over 1 epoch during the SFT stage and on 47.3K preference data for 1 epoch during the DPO stage.
Results As shown in Table 1, on the FrameCapEval benchmark, ProgressCaptioner greatly outperforms existing open-source VLMs of similar capacity and even the (much larger) latest Gemini-1.5-Pro and GPT-4o. We observe that strong language-backed VLMs like GPT-4o show high caption matching accuracy, whereas Qwen2-VL excels in progression detection, reducing hallucination. However, it tends to produce less detailed captions, leading to lower caption matching accuracy. In contrast, ProgressCaptioner effectively balances precision and detail in frame-wise captioning, consistently leading the benchmark across both in-domain and out-of-domain datasets.
Figure 5 provides qualitative comparisons on three action sequences. Consider the (a) bowling sequence for instance: baseline models erroneously suggest progression in frame 2, like “arm forward”, exemplifying the common issue of temporal hallucination in current VLMs. This issue recurs in the other two sequences. Conversely, ProgressCaptioner delivers high-quality captions that precisely characterize action progress in each frame. See Supp. for more qualitative examples and an ablation of ProgressCaptioner.
User Study Figure 6 presents the user study results, where ProgressCaptioner is compared against four of the strongest competitors: LLAVA-OV, Qwen2-VL, Gemini-1.5-Pro and GPT-4o. Each participant is presented with five captions produced by these models and is tasked with selecting the top-2, with an additional “none” option if the captions are deemed inadequate. ProgressCaptioner emerges as the most preferred model, with an average best caption selection rate of 31.6%—2$\times$ to 3.6$\times$ better than the comparably sized best models from the literature [62, 33], and even surpassing top-tier proprietary models that enjoy significant scaling advantages. While our model outperforms all open-source and proprietary models for top-1 preference, the more forgiving top-2 metric brings the proprietary closed models back in the game, though our model remains competitive even there (50.3% for GPT-4o vs. 47.3% for ours). These findings underscore our model’s strong ability to produce accurate, temporally fine-grained captions.
4.2 Applications of Video Frame Captioning
ProgressCaptioner offers progress-aware frame-wise captions, which hold great potential for many real-world applications. We explore several practical use-cases below.
Keyframe Selection Our first use-case leverages ProgressCaptioner’s temporally precise captions as an intermediate representation to identify keyframes within a densely sampled sequence, aided by an LLM (see Supp for details). Figure 7 provides two examples, showcasing how ProgressCaptioner’s produced captions allow selecting distinct frames that effectively capture different stages of the welding and squeezing grapefruit action. While recent video summarization work [24] explores using VLMs and LLMs for keyframe selection, it aims to identify coarse-grained events within long videos, which is not adequate for our problem scenario. See Supp. for a side-to-side comparison and more qualitatives that underscore this distinction.
Keyframes for Action Recognition Not only is keyframe selection useful for human viewers to quickly preview a longer video, but it can also extract the most informative portions of a video to benefit activity recognition [29]. To illustrate, we next apply our keyframe selection mechanism to Kinetics [7] Temporal [57] subset that necessitates multi-frame reasoning. Given that the original video clips are short (sampled at 1FPS, resulting in sequences of 10 frames), we employ two models that take four frames as input: Slow backbone from SlowFast [17] and X3D-XS [16]. We prompt GPT-4o to select four representative frames from frame-wise captions produced by ProgressCaptioner. We take the two model checkpoints that have been trained on the Kinetics training set and replace uniformly sampled frames with our selected keyframes during inference. Figure 8 presents a qualitative comparison, highlighting performance gains such as a +1.7% increase in top-1 accuracy for both SlowFast and X3D models. Even among just 10 candidate frames, our method’s fine-grained ability to identify the 4 most informative ones translates into better recognition.
Figure 8: On Kinetics test videos, ProgressCaptioner selects four frames that are more informative of the action than uniform sampling, resulting in improved action recognition accuracy.
Advancing Video Understanding The precise, frame-wise captions generated by ProgressCaptioner enhance our understanding of videos. To demonstrate this, we consider two video tasks that demand temporally fine-grained understanding: (1) frame-wise classification on HowToChange [72] and Penn Action [81], and (2) video question answering (QA) on NExT-QA [69] (ATP-Hard [5]). These tasks are chosen because they challenge the model to comprehend not just the overarching content of a video, but also the more fine-grained event progression within a video. The HowToChange and Penn Action test sets provide frame-wise labels detailing object state changes or action phases, requiring frame-level understanding. Similarly, NexTQA (ATP-Hard) poses temporally challenging questions that demand multi-frame reasoning, such as determining event order, emphasizing the need for precise temporal comprehension. For baseline comparisons, we evaluate against the LLAVA-OV-7B [33] model, from which ProgressCaptioner is initialized, to highlight the enhancements that our specialized training on FrameCap brings to video understanding tasks. For the first task, as we pioneer a zero-shot, language-guided approach to this traditionally vision-centric problem (details below), no other zero-shot baselines exist. For the second task, we compare ProgressCaptioner against two existing zero-shot approaches [45, 64].
Zero-shot Frame Classification We repurpose zero-shot frame-wise classification task into a multi-choice QA format, using frame-wise captions to guide an LLM in identifying the correct label per frame, evaluating caption accuracy and granularity (Figure 9 left). Results (Figure 9 right) show that ProgressCaptioner consistently outperforms LLAVA-OV across both datasets. Notably, our training involves no signals related to these frame-wise labels, underscoring its generalizability and effectiveness in enhancing video frame-level understanding.
Figure 9: ProgressCaptioner delivers precise and detailed per-frame descriptions, leading to enhanced zero-shot frame-wise classification performance when compared with LLAVA-OV.
Video QA Finally, we report results using frame-wise descriptions for video QA, where an LLM (we use GPT-4o) is employed to answer questions on NExT-QA (ATP-Hard) set. As shown in Table 2, ProgressCaptioner achieves the best results on this benchmark, outperforming the previous leader VideoAgent [64] by +3.4%. Compared with a similar setup using LLAVA-OV, ProgressCaptioner achieves a +4.7% gain in the temporal subset, highlighting its superior ability to produce fine-grained, temporally precise descriptions and bring enhanced video understanding.
 Model Acc@C Acc@T Acc@All
VFC [45] 32.2 30.0 31.4
VideoAgent [64] 57.8 58.8 58.4
LLAVA-OV [33] + GPT-4o 62.6 53.4 58.8
ProgressCaptioner + GPT-4o (ours) 64.4 58.1 61.8
Table 2: Video QA results on NExT-QA (ATP-Hard). C and T denote causal and temporal subsets, respectively.
5 Conclusion
We introduce progress-aware video frame captioning, which necessitates a significant enhancement in current captioning models’ capability to describe temporal action progression precisely. Towards this end, we develop ProgressCaptioner and show its effectiveness in enhancing the temporal precision and alignment of captions with corresponding frames. Furthermore, we demonstrate its practical applications: keyframe selection and enhanced video understanding. By setting a new standard for temporal precision in video captioning, we hope our work inspires further development in this evolving domain.
Supplementary Material
1 Dataset
1.1 FrameCap Training Data
To construct the FramePair dataset, we employ a suite of open-source VLMs as captioners for initial pseudo label generation, including VILA [40], Qwen2-VL [62], LLAVA-Next-Video [35], LLAVA-Video [82] and LLAVA-OV [33]. Training videos are sourced from HowToChange and COIN, with frames extracted at 1FPS. We prepare pairs of frames for stage-I and multi-frame sequences for stage-II; the frame sequence length ranges from 3 to 6, as our preliminary experiments suggest that extending beyond 6 frames causes multiple issues with our captioners, such as overly brief captions, memory overflows, and great temporal mismatches.
We then process the data through our custom-designed tasks: progression detection and caption matching, to filter for high-quality data. The progression detection uses LLAMA-3.1-70B-Instruct [14], and for caption matching, we use VILA [40], chosen for its open-source availability and strong performance. Specifically, we assess caption matching precision by comparing model-generated answers against human responses on a subset of 90 questions. Gemini-1.5-Pro [53] achieves a precision of 0.89, while VILA achieves 0.75, the highest among open-source VLMs. Given that Gemini-1.5-Pro API usage incurs a cost, we reserve it for evaluation while utilizing the cost-free VILA as the caption matching evaluation VLM during the pseudo labeling stage.
For each frame sequence, the caption sequence that passes and fails these evaluations forms our preference data, which is utilized for DPO training of ProgressCaptioner. See Figure 10 and  11 for examples of frame pair data obtained from progression detection and caption matching, respectively, and Figure 12 for an illustration of the frame sequence data preparation process. Table 3 provides a summary of the training data statistics. The first data preparation stage collects a total of 240K frame-caption pairs for supervised fine-tuning (SFT) and 21K preference pairs for direct preference optimization (DPO). The second stage further expands the dataset to include 34K multi-frame and caption sequences for SFT, along with 26K frame-caption sequences for DPO.
Figure 10: Example of a frame pair (decided by progression detection). The upper caption pair is marked as “accepted” by the evaluation LLM, aligning with the ground truth progression label (no progression), while the lower caption pair is marked as “rejected” because it incorrectly suggests progression.
Figure 11: Example of a frame pair (decided by caption matching). The upper caption pair is marked as “accepted” since the evaluation VLM correctly answers the caption matching questions as (A, B), demonstrating good alignment. In contrast, the lower pair is “rejected” due to its answers (B, A), indicating poor correspondence between the frame and the generated captions.
Figure 12: Example of a frame sequence. Progression detection is first applied to each adjacent frame pair to determine the visual-change label and identify $M$ distinct frames. Caption matching then evaluates the captions corresponding to these $M$ frames. The upper caption sequence is “accepted” as the evaluation VLM correctly answers (A, B, C, D), whereas the lower caption sequence, leading to erroneous responses, is marked as “rejected”.
 Dataset # Videos # Frames # Pair # Seq
SFT DPO SFT DPO
HowToChange [72] 7,812 101,369 83,383 8,453 13,602 8,362
COIN [59] 9,030 103,791 156,858 12,622 20,704 17,826
Total 16,842 205,160 240,241 21,075 34,306 26,188
Table 3: We propose the FrameCap data collection, offering large-scale frame and caption sequences for fine-grained frame-level video captioning.
1.2 FrameCapEval Benchmark
For the FrameCapEval benchmark, we source videos from four action-focused datasets: HowToChange [72], COIN [59], Penn Action [81] and Kinetics [7]. We ensure a balanced selection of videos from each action category across these datasets and follow their original validation or test splits. We are mindful of the single frame bias [32]—a recognized issue in video understanding where some actions are not distinctly temporal and can be adequately depicted with a single frame. To address this, we conduct a manual verification of all videos to eliminate frame sequences that lack fine-grained action progression, as these scenarios are straightforward and can be adequately managed by image captioning models. Frames are extracted at 1 FPS and grouped using $K$-means clustering based on CLIP features [51], with $K$ determined by silhouette scores [55] and ranging from 3 to 6. To each sequence, we add a frame with the smallest CLIP feature distance from a randomly chosen frame, so that the final sequence captures scenarios with and without action progression. See Table 4 for detailed evaluation data statistics.
 Dataset # Videos # Frames
HowToChange [72] 306 (102) 1101
COIN [59] 271 (139) 1063
Penn Action [81] 51 (47) 235
Kinetics600 [7] 56 (52) 451
Table 4: FrameCapEval data statistics. The numbers in parentheses represent the count of videos used for caption matching. We manually verify all selected frame sequences to assign action progression labels and filter out low-quality (easy) examples lacking clear action progression. This process ensures a robust testbed for evaluating a model’s capability to generate temporally fine-grained descriptions.
FrameCap and FrameCapEval offer unique resources for temporally fine-grained descriptions at the frame level, which can be a valuable enhancement to current VLM’s training data. We will publicly release the two datasets and hope that these resources help advance the temporal precision in video understanding capabilities of VLMs.
2 Experiments
2.1 Experimental setup
Evaluation Metric Design
Progression detection evaluates a model’s action progress awareness, using caption pairs generated for each frame pair. It functions as a binary classification task, where label = 0 identifies scenarios with no visual progression to detect hallucinations, and label = 1 signifies visual progression to assess the model’s ability to capture detailed temporal changes. We measure performance using balanced accuracy, which averages the true positive and true negative rates to account for data imbalance. To enhance the reliability and quality of our evaluations, we manually annotate visual progression between frames in the FrameCapEval dataset. Llama-3.1-70B-Instruct [14] is employed as the evaluation LLM to determine if a caption pair describes visual progression.
Caption matching assesses both the accuracy and the temporal granularity of captions. The evaluation is conducted on $T$-frame sequences that depict action progression, which are manually validated to ensure reliability. Gemini-1.5-Pro [53] is adopted as the evaluation VLM and tasked with performing the frame-wise caption matching task. We measure sequence-level accuracy, defined as the proportion of sequences where every frame is correctly identified by the evaluation VLM among all test sequences. It reflects how many caption sequences are entirely correct, which effectively rules out the possibility of random guessing being successful for a few frames within the sequence, providing a more robust assessment of caption sequence quality.
For the user study, 15 participants reviewed 85 frame sequences (364 frames) in total, randomly sampled from the FrameEvalBenchmark. We evaluate the captions from four leading models—two open-source (LLAVA-OV [33], Qwen2-VL [62]) and two proprietary (Gemini-1.5-Pro [53], GPT-4o [2])—alongside our ProgressCaptioner. Note that image captioning baselines are excluded due to their excessively lengthy captions and complete lack of temporal coherence. Participants are presented with captions produced by these five models, randomly shuffled for each sequence, and asked to choose the best and second best (with an additional “none” option available) for each frame’s caption. The average selection rate per model is reported, providing insights into subjective caption quality preferences.
Implementation
The Stage-I (frame pair captioning) and Stage-II (frame sequence captioning) models are trained with the same hyperparameters and undergo the same training processes: SFT followed by DPO. In the SFT phase, learning rates are set at 1e-5 for the LLM and projector, and 2e-6 for the vision encoder, with a batch size of 64. For DPO, the learning rate is reduced to 5e-7 with a batch size of 8. We set the preference scaling parameter $\alpha$ = 1.0 and the temperature parameter $\beta$ = 0.2.
During inference, ProgressCaptioner takes frame sequences ranging from 2 to 6 frames. This limit is set because, as discussed earlier, all models experience severe performance degradation with longer frame sequences; hence, we cap at 6 frames when preparing training data and keep the inference protocol consistent with training. For sequences exceeding this length, ProgressCaptioner can operate in a sliding window mode.
For results in Section 4.1, direct inference is applied on $T$ frames. For results in Section 4.2, we employ a 2-frame sliding window, where ProgressCaptioner performs frame pair captioning (except for NeXT-QA, where we uniformly sample 6 frames from the original video and apply direct inference on these 6-frame sequences without a sliding window). A single frame ($v_{t}$) can receive two captions: one from the pair ($v_{t-1}$, $v_{t}$) and another from ($v_{t}$, $v_{t+1}$). We concatenate the two captions for frame classification tasks to provide richer contextual information, aiding the LLM in frame label prediction. For keyframe selection, we use the caption from the pair ($v_{t-1}$, $v_{t}$) for $v_{t}$ to maintain caption sequence coherence.
2.2 Prompt used
We design the following prompt for VLMs to perform the frame-wise video captioning task:
Caption Generation Prompt Instructions: These are T frames extracted from a video sequence depicting action. Provide a detailed description for each frame. Requirement: (1) Ensure each frame’s description is specific to the corresponding frame, not referencing other frames. (2) The description should focus on the specific action being performed, capturing the progression of the action. There is no need to comment on other elements, such as the background or unrelated objects. Reply with the following format: <Frame 1>: Your description <Frame T>: Your description.
where T represents the number of frames in the sequence, and action is the video-level action label. The prompt is selected based on preliminary experiments on a small set of data, and we manually review the generated captions to ensure their effectiveness. We use the same prompt consistently for pseudo labeling training data and for evaluating current VLMs, both for our model and existing ones.
The progression detection prompt provided to the LLM is as follows:
Progression Detection Prompt (Pseudo-labeling) Instructions: You will be provided with two image descriptions. Your task is to determine the relationship between the two images based on these descriptions. Image 1 description: desc1 Image 2 description: desc2 Choose the most appropriate option from the following: A. The images likely look similar (no significant change). B. There are noticeable changes between Image 1 and Image 2. C. It is not possible to determine the similarity or difference based on the descriptions.
Progression Detection Prompt (Evaluation) Instructions: You will be provided with two image descriptions depicting an action. Your task is to determine the relationship between the actions in the two images based on the descriptions provided. Action: action The image descriptions are: Image 1: desc1 Image 2: desc2 Choose one of the following options: A. Action Progression: The action has advanced from Image 1 to Image 2 (e.g., more of the task has been completed in Image 2). B. No Action Progression: The action remains the same between Image 1 and Image 2 (e.g., the images may show a change in viewpoint, hand position, or slight object adjustments, but the action itself has not progressed). C. Uncertain: It is unclear whether the action has progressed or not.
In these prompts, desc1 and desc2 represent the descriptions of Image 1 and Image 2, respectively, and action is the video-level action label. The progression detection prompts differ between training and evaluation as they serve distinct purposes. For training, we aim to identify visually different frames within a sequence to ensure that the frame sequences processed later by caption matching are composed of distinct frames. Therefore, the training prompt focuses on detecting any visual changes, regardless of their nature. For evaluation, the objective shifts to determining whether the caption sequence is progress-aware; we manually annotate each frame sequence with progression labels for this purpose. As such, the evaluation prompt is designed to discern whether there is action progression or no action progression, rather than identifying simple visual changes. It is important to note that “changes” can encompass broader aspects than “progression”, as explained in the prompt, changes may include viewpoint change or background object adjustments, which do not necessarily indicate a progression in the ongoing action.
Consider a sequence of $M$ visually distinct frames $\mathcal{V}_{M}=\{v_{i}\}_{i=1}^{M}$, as detailed in Sec. 3.2. We task an evaluation VLM to perform caption matching for each frame $v_{m}\in\mathcal{V}_{M}$ with the following prompt:
Caption Matching Prompt Which caption best describes the image? <Frame $v_{m}$> A. Caption $\hat{c}_{1}$vdots M. Caption $\hat{c}_{M}$ M+1. None of the above descriptions match the image, are hard to determine, or contain incorrect information about the image. Reply with only the corresponding letter (A, B, C, etc.)
where <Frame $v_{m}$> denotes the image input (the $m$-th frame), and $\{\hat{c}_{i}\}_{i=1}^{M}$ is the caption sequence to be evaluated.
2.3 Results
Ablation Study
Table 5 presents an ablation study using HowToChange videos from FrameCapEval, focusing on three key variables: (1) comparisons between Stage-I and Stage-II models; (2) the effect of training datasets—HowToChange alone versus HowToChange combined with COIN; (3) the impact of SFT alone versus SFT followed by DPO. The results demonstrate that all three factors are crucial for optimal performance. First, the Stage-I model, limited to frame pair captioning, does not provide caption matching accuracy for $T$-frame sequences and shows lower progression detection performance compared to the Stage-II model, which benefits from additional frame sequence training (see row 1 vs. row 2). Second, regarding training data, while evaluation is conducted on HowToChange, incorporating COIN data for training greatly improves performance, particularly in caption matching, highlighting the benefits of data scaling (see row 2 vs. row 4). This indicates potential for further enhancements by adding more datasets in the future. Finally, direct preference optimization (DPO) proves critical as its absence leads to great performance declines (see row 3 vs. row 4).
 Model Training Data Training Strategy Cap Prog Stage-I HTC SFT+DPO 70.6 Stage-II HTC SFT+DPO 28.4 73.1 Stage-II HTC+COIN SFT 24.5 68.3 Stage-II HTC+COIN SFT+DPO 37.3 73.6   Table 5: Ablation study of ProgressCaptioner on HowToChange (HTC) evaluation videos: examining the impact of model stages, training data, and optimization strategies on performance. More Qualitative Results Supplementing Figure 5 in the main paper, Figures 1315 provide more qualitative predictions, comparing ProgressCaptioner with the four leading VLMs. As can be seen from these examples, while all baseline VLMs exhibit temporal inaccuracies in their descriptions, ProgressCaptioner consistently provides temporally precise and progress-aware captions, highlighting its superior performance.
Figure 13: Qualitative comparisons of ProgressCaptioner with SOTA VLMs (I). Red text identifies inaccuracies in the generated captions, while blue text highlights how our progress-aware captions build on prior content to clearly delineate what is changing or continuing.
Figure 14: Qualitative comparisons of ProgressCaptioner with SOTA VLMs (II). Red text identifies inaccuracies in the generated captions, while blue text highlights how our progress-aware captions build on prior content to clearly delineate what is changing or continuing.
Figure 15: Qualitative comparisons of ProgressCaptioner with SOTA VLMs (III). Red text identifies inaccuracies in the generated captions, while blue text highlights how our progress-aware captions build on prior content to clearly delineate what is changing or continuing.
Figure 16: Captions produced by ProgressCaptioner and processed by an LLM enable us to automatically select representative frames that clearly depict action progression from densely sampled frame sequences. For each frame sequence, the bottom left box displays discarded frames alongside their captions, while the bottom right box showcases selected frames and their corresponding captions. This process effectively removes duplicate frames that depict the same action progression and enhances the selected frames with captions.
Keyframe Selection
We propose to utilize frame-wise captions from ProgressCaptioner to select frames that depict action progression. The key idea is to “encode” a sequence of densely sampled video frames into per-frame captions, allowing an LLM to subsequently “decode” and identify key frames from this rich textual representation. The temporally fine-grained descriptions act as a condensed frame representation, focusing on action progression while remaining robust to visual disturbances such as changes in viewpoint or background objects. Figure 16 illustrates one potential design for such a keyframe selection feature. With ProgressCaptioner, we employ a sliding two-frame window for captioning, followed by an LLM (we use Llama-3.1-70B-Instruct) processing the generated captions. Specifically, for a sequence of densely sampled frames $\{v_{t}\}_{t=1}^{T}$, starting from $t=1$, ProgressCaptioner generates caption ($c_{1}$, $c_{2}$) for ($v_{1}$, $v_{2}$). We then ask the LLM to determine if there is action progression between $c_{1}$ and $c_{2}$. If the answer is yes, frame $v_{2}$ gets selected; if no, $v_{2}$ is skipped to avoid redundancy as it likely depicts the same action stage as $v_{1}$. The process is repeated by advancing the window to ($v_{2}$, $v_{3}$) and continuing through the sequence.
Our approach offers two key advantages: (1) it efficiently filters out non-essential frames to ensure that selected frames distinctly represent action progression, and (2) it dynamically determines the size of the keyframe set based on the sequence content, eliminating the need for manually specifying the number of frames to sub-sample. To better illustrate this, we compare our method with the pseudo labeling strategy used in a recent video summarization work, V2Xum [24]. V2Xum employs an image captioning model followed by an LLM to perform extractive document summarization based on per-frame captions for keyframe selection.
As shown in Figure 17, V2Xum’s approach results in duplicate keyframes for sequence (a), where the first and second frames depict the same action progression despite a viewpoint change, and the last three frames similarly represent the action progression of oranges being sliced in half. In contrast, our method, leveraging the more accurate and temporally fine-grained captions produced by ProgressCaptioner, precisely identifies three distinct stages of this slicing action sequence. For sequence (c), V2Xum selects only one frame from the first four, despite depicting various stages of cutting a sausage (from whole to partially cut, fully cut, and then to chunks). Conversely, our approach accurately identifies all these frames as markers of action progression. It adaptively determines the size of the keyframe set, which can vary from small to large depending on the actual content, offering flexibility without requiring manual specification.
To conclude, our keyframe selection approach effectively highlights critical moments within action sequences. We believe such a system has significant potential for providing focused insights in educational tutorials and sports analysis, benefiting learners and analysts alike.
Figure 17: Comparison of our keyframe selection with V2Xum [24]. Leveraging precise and progress-aware captions from ProgressCaptioner, our approach selects keyframes that accurately represent stages of the action process. In contrast, V2Xum’s method often includes duplicate frames or overlooks frames that show subtle but important differences.
Limitations
Despite the enhanced performance of ProgressCaptioner, it still faces several challenges. Firstly, while we have developed an advanced pseudo labeling refinement process, the training data sourced from existing VLMs inherently limits the quality of the captions. Moreover, the automation of data filtering using evaluation LLMs and VLMs introduces noise—though less costly, it’s not as reliable as human annotation. Secondly, we observe that captioning longer frame sequences presents increased difficulties; for instance, accurately captioning six-frame sequences is notably more challenging than two-frame sequences. Addressing this challenge to extend ProgressCaptioner’s capabilities to handle longer sequences remains a critical area for future development. In addition, Figure 18 illustrates a failure case where ProgressCaptioner produces identical captions for the last four frames, failing to recognize fine-grained spatial changes—an area that current VLMs consistently fall short of. This underscores the need for further advancements in this area.
Finally, we emphasize that the task of video frame captioning introduces a significant challenge by demanding high temporal precision. We recognize the limitations of ProgressCaptioner in its current stage and view this work as an initial step toward resolving this problem.
Figure 18: One failure case of ProgressCaptioner, where it fails to discern fine-grained spatial differences among the last four frames and thus produces identical captions.